OpenStudy (anonymous):

. Caroline purchases $75.00 worth of clothing. If the sales tax is 7%, what is the total cost? 1 a. $5.25 b. $69.75 c. $80.25 d. $81.00

OpenStudy (haleyelizabeth2017):

75*1.07.....try that :)

OpenStudy (anonymous):

8025.?

OpenStudy (jskhupmang):

why is it 1.07? shouldn't it be .07?

OpenStudy (haleyelizabeth2017):

No because it is sales tax so you add it to the total cost...it's the same as doing 75*0.07 and then adding that to 75, just a step easier ;)

OpenStudy (anonymous):

75 * 1.07 = 80.25

OpenStudy (haleyelizabeth2017):

not that much lol 1.07*75 is indeed 80.25

OpenStudy (jskhupmang):

75*.07 = x x+75 is the total ?

OpenStudy (haleyelizabeth2017):

Yep :)

OpenStudy (haleyelizabeth2017):

Which is the same as 1.07*75, just this is a simpler way ;)

OpenStudy (anonymous):

is that it?

OpenStudy (haleyelizabeth2017):

80.25 is indeed the answer :D

OpenStudy (jskhupmang):

how did you just know 1 infront of .o7?

OpenStudy (haleyelizabeth2017):

because 1* something = the original number. then if you do 1.07, that adds the tax, or the extra cost :)

OpenStudy (jskhupmang):

ok thxs

OpenStudy (anonymous):

75*100%+75*7%=75*107%=75*1.07=80.25

OpenStudy (haleyelizabeth2017):

You're welcome :)

OpenStudy (haleyelizabeth2017):

Good way of explaining it @robtobey :)

OpenStudy (anonymous):

Glad it helped you out.

OpenStudy (haleyelizabeth2017):

Eh?

OpenStudy (haleyelizabeth2017):

I already knew how to do it lol just didn't know how to explain it right :P but good job anyways teehee